Determine the number of solutions and classify the type of solutions for each of the following equations. Justify your answer.
4x2 + 16x + 16 = 0

Answer by RAY100(1637)

4x^2 + 16x +16 =0
,
4 (x^2 + 4x + 4) =0
,
check Discriminant, remember, D less than 0, no solutions
,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,D =0,,,,,one solution
,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,D > 0,,,,,two solutions
,
a=1,,b=4,,,c=4,,,,and D = sqrt ( b^2 - 4 * a * c ) = sqrt ( 16 -4 * 1 * 4 ) =0
,
therefore one solution
,
factoring ( x^2 +4x +4 ) = (x+2) (x+2 ),,, x=-2,,and -2,,
,
this is a duplicate solution and counts as 1 (one) solution
,
The function is a positive parabola, cup pointed up, with vertex just touching y axis, at x=-2